7) Factorizar 3x2 + 9x - 30
Answer #1

3x +92-30=3x²+158-62-30 -: 154-62=93) = 3x(x+5) - 6(x+5) (24+5)(3x-6) Thus 32 +99-30 = (x+5)(34-6)